OpenStudy (anonymous):

Write an equation for a line in slope-intercept form that passes through the points (3, 4) and (1, -2). y = x – 1 y = x + 1 y = 3x – 5 y = 3x + 5

OpenStudy (anonymous):

its y=3x-5

OpenStudy (anonymous):

m = (y2-y1)/(x2-x1), and you have two ordered pairs (3, 4) and (1, -2) to plug in values from. Solving yields m = 3 to find the equation for a line (in slope-intercept form), you may use: y = m(x-x0) +y0 plug in values to get y=3(x-3) + 4 which simplifies to option 3 y=3x-5